# Pressure-temperature phase diagram calculations using polynomial machine learning potentials: A comprehensive study based on global structure prediction and self-consistent phonon calculations

### v1

@@ I Introduction
Many elemental systems and compounds are known to exhibit polymorphism, wherein they can crystallize into multiple distinct crystal structures. Notably, numerous polymorphs have been observed, especially under high-pressure conditions. For instance, experimental studies on elemental silicon (Si) have identified seven distinct phases that exist in sequence with increasing pressure: Si-I (diamond), Si-II ($\beta$-Sn), Si-XI (space group $Imma$), Si-V (simple hexagonal, SH), Si-VI (space group $Cmce$), Si-VII (hexagonal close-packed, HCP), and Si-X (face-centered cubic, FCC) within the pressure range of 0–100 GPa at room temperature [1, 2, 3]. To predict such crystal structures under varying pressure conditions, ranging from ambient to high pressures, global crystal structure searches using density functional theory (DFT) calculations have proven to be a powerful and effective tool [4, 5, 6, 7].
@@
Machine learning potentials (MLPs) have been widely used to perform large-scale simulations and a vast number of systematic calculations accurately and efficiently, which would be computationally expensive using DFT calculations [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. These MLPs represent interatomic interactions by utilizing a range of structural features that describe the neighboring atomic distribution, in conjunction with machine learning models such as artificial neural networks, Gaussian process models, and linear models. Typically, MLPs are developed from extensive datasets derived from DFT calculations, which enables them to outperform conventional interatomic potentials in predicting various material properties. Recent studies have shown that MLPs significantly accelerate global structure searches by facilitating efficient energy and force calculations during local geometry optimizations [33, 34, 35, 36, 37, 38, 39]. This acceleration allows for robust global structure searches across a wide search space, significantly reducing the likelihood of missing the true global minimum structures. The MLPs are also useful for performing reliable global structure searches under high-pressure conditions. Meanwhile, MLPs capable of accurately predicting a wide variety of structures under high-pressure conditions, including numerous hypothetical structures, are essential, and systematic global structure searches must be conducted.
Identifying globally stable structures at finite temperatures is also a computationally demanding task. To estimate phase stability at finite temperatures among local minimum structures obtained from global structure searches, self-consistent phonon approaches [40, 41, 42, 43, 44] can be employed with the effective use of MLPs. These approaches account for both harmonic and anharmonic vibrational effects and predict temperature-induced dynamical stability in structures that are dynamically unstable at lower temperatures. However, even when using MLPs, performing self-consistent phonon calculations in conjunction with global structure searches remains challenging for the following reasons. First, a comprehensive enumeration of structures with energy values close to the global minimum is necessary. Second, systematic free energy calculations using self-consistent phonon approaches must be performed for the enumerated local minimum structures. Such calculations are especially demanding for structures with fewer symmetry operations because they involve a large number of independent force-constant components, making the estimation of effective force constants computationally expensive. Furthermore, MLPs that can accurately perform self-consistent phonon calculations across various local minimum structures must be developed. Consequently, while MLPs are expected to be highly effective for identifying globally stable structures at finite temperatures, a comprehensive methodology combining global structure searches and free energy evaluations remains indispensable.
This study presents an efficient and robust approach that utilizes MLPs for crystal structure predictions under various pressure and temperature conditions. The performance of the proposed procedure is evaluated using elemental Si, which exhibits several experimentally observed polymorphs. To comprehensively and globally enumerate local minimum structures, this study employs an iterative procedure that combines random structure search (RSS) with the development of an MLP, which demonstrates high predictive power across a wide range of structures. This methodology has been previously applied in global structure searches using the polynomial MLP [37, 38, 39]. Subsequently, the stochastic self-consistent harmonic approximation (SSCHA) [45, 46, 47] is employed to investigate the phase stability among the local minimum structures identified through the global structure search. The polynomial MLP is used to perform the calculations required for this study. The polynomial MLP is constructed based on polynomial rotational invariants systematically derived from order parameters expressed in terms of radial and spherical harmonic functions. Polynomial MLPs developed for general purposes have demonstrated high predictive power across a wide range of crystalline [48, 49, 50, 51] and liquid states [52].
An overview of the procedure proposed in this study is provided in Fig. 1, and the paper is organized as follows. Section II describes the formulation and construction procedure of the polynomial MLP. Section II.5 presents a grid search approach for obtaining an optimal MLP, which is employed in the global structure search and phase stability analysis. Section III details the iterative procedure for a robust global structure search and presents the globally stable and metastable structures identified by the search. Section IV introduces the phase stability analysis procedure and provides the pressure-temperature phase diagram calculated using the proposed methodology. Section IV.4 assesses both the computational requirements and the predictive accuracy of the MLP for SSCHA calculations. Finally, Sec. V concludes the study.
@@ II.1 Formulation of the polynomial MLP
In addition to the model given by Eq. (5), simpler models composed of a linear polynomial of structural features and a polynomial of a subset of the structural features are also introduced, such as
@@ II.2 Hybrid polynomial MLP models
In this study, we propose an extended polynomial MLP model, which is represented as the sum of multiple polynomial MLP models using different sets of input parameters. Such a model will be referred to as a hybrid polynomial MLP model. The atomic energy in the hybrid polynomial MLP is expressed as the sum of two distinct MLP models, $E^{(i)}_{\rm{mdl1}}$ and $E^{(i)}_{\rm{mdl2}}$, given by
@@
Here, $D^{(i)}_{\rm{mdl1}}$ and $D^{(i)}_{\rm{mdl2}}$ represent sets of structural features generated using different input parameter sets, such as the cutoff radius, the number of radial functions, and the truncation of polynomial invariants. Specifically, these parameters include the maximum angular momentum of spherical harmonics and the polynomial order of the invariants. These hybrid polynomial MLP models extend the framework presented in Eq. (5) and are expected to efficiently include a broader range of structural features derived from various input parameters, without requiring the consideration of intersection polynomials between the individual models.
@@ II.3 Datasets
We begin the development of MLPs using a dataset from DFT calculations, comprising approximately 12,000 structures for elemental Si, which was previously developed in other studies [53, 51]. The procedure used to generate this dataset is summarized as follows. First, the atomic positions and lattice constants of 86 prototype structures [48] were optimized using DFT calculations under zero pressure. These prototypes consist of single elements with zero oxidation state from the Inorganic Crystal Structure Database (ICSD) [59], including metallic close-packed structures, covalent structures, layered structures, and those reported as high-pressure phases. Next, supercells of these optimized structures were used to introduce random lattice expansions, lattice distortions, and atomic displacements. DFT calculations were finally performed on these supercells to generate the dataset. This dataset will be referred to as dataset 1.
Dataset 1 is derived solely from prototype structures optimized at zero pressure and therefore includes only a limited number of structures with small volumes. As a result, the MLP trained on this dataset has relatively low predictive power for various properties under high-pressure conditions, as demonstrated later. To develop MLPs with higher predictive accuracy under high-pressure conditions, we generate additional structures from prototype structures optimized at elevated pressures using the following procedure.
First, the 86 prototype structures used to construct dataset 1 are optimized using DFT calculations at 25, 50, 75, and 100 GPa. Next, supercells of these optimized prototypes are used to create candidate structures by randomly introducing lattice expansions, distortions, and atomic displacements, following the same procedure as dataset 1 [53]. A parameter controlling the upper bounds of the magnitude for these transformations is set to 0.3 $\rm{\AA}$. Finally, to reduce uncertainty more efficiently than random selection, we apply the structure-selection procedure described in Ref. [48], extracting a total of 4,000 structures from the candidate set. The dataset comprising these structures, combined with dataset 1, will be referred to as dataset 2. When the dataset is used to estimate MLPs through regression, it is randomly split into training and test sets, with 90 % allocated to training and 10 % to testing.
@@ II.4 MLP estimation
Here, $\lambda$ represents the regularization parameter optimized to minimize prediction errors on the test set, and $\bm{W}$ is a diagonal matrix whose entries weight the energy, force, and stress tensor data. The weight settings for the energy and force entries are consistent with those proposed in Ref. [38]. In these settings, larger weights are assigned to data entries with lower energy values and smaller forces, which is critical for accurately modeling structures near local minima. To preserve high predictive accuracy of the regression model for the energy, force, and stress tensor data, the stress tensor entries are assigned relatively smaller weights than the energy entries.
II.5 MLP optimization
@@
As demonstrated in Ref. [48], the accuracy and computational efficiency of the polynomial MLP are trade-off properties. We therefore obtain a set of Pareto-optimal MLPs with different trade-offs between these properties through a grid search. This search covers 523 single polynomial MLP models and 640 hybrid polynomial MLP models. In the single polynomial MLP models, the cutoff radius ranges from 4 to 12 ${\rm{\AA}}$, and polynomial invariants are considered up to approximately 25,000. In the hybrid polynomial MLP models, two separate models are combined. The first model uses smaller cutoff radii (4, 5, or 6 ${\rm{\AA}}$) and 3,000–25,000 polynomial invariants, capturing a wide range of coordination environments. The second model adopts larger cutoff radii (6, 7, 8, 10, or 12 ${\rm{\AA}}$) and a varying number of polynomial invariants (1,000 to 25,000). These polynomial MLPs are developed from datasets using the pypolymlp code [53, 66].
Figure 3: Prediction error and computational efficiency of polynomial MLPs obtained from the grid search in elemental Si. The blue and green open circles represent the properties of MLPs constructed using single polynomial MLP models and hybrid polynomial MLP models, respectively. Computational efficiency is assessed based on the computation time using a single core of the ${}^{\text{\textregistered}}$ ${}^{\text{\textregistered}}$ E5-2630 v3 (2.40 GHz). The orange open squares denote the Pareto-optimal MLPs across both single and hybrid polynomial MLP models, while the blue open diamonds represent the Pareto-optimal MLPs from the single polynomial MLP models only. The red closed square indicates the selected MLP for conducting the global structure search and phase stability analysis.
Figure 3 shows the prediction error and computational efficiency of the polynomial MLPs obtained through the grid search for elemental Si. The prediction errors are evaluated using the root mean square errors (RMSEs) of the energy. To ensure practical relevance, these RMSEs are calculated using the test datasets that exclude structures with exceptionally high energies or large forces. The computational efficiency is assessed by measuring the elapsed time to compute the energy, forces, and stress tensors 10 times for a structure containing 1000 atoms. As the model complexity of the polynomial MLP increases, the prediction error generally decreases. Among the Pareto-optimal MLPs, the polynomial MLPs constructed using hybrid polynomial MLP models demonstrate lower prediction errors than those constructed with single polynomial MLP models when the computational cost exceeds 1.5 ms/step/atom. Hence, the hybrid polynomial MLPs can provide Pareto-optimal MLPs with a better trade-off between prediction error and computational cost.
@@
A reasonable MLP with a computational time of 3.1 ms/atom/step is selected from the set of Pareto-optimal MLPs, balancing prediction accuracy and efficiency. It is designated as the optimal MLP for performing global structure searches and phase stability analyses, as indicated by the red closed square in Fig. 3. Its RMSEs for energy, force, and stress tensor are 2.8 meV/atom, 0.055 eV/Å, and 35.9 meV/atom, respectively. Figure 4(a) compares cohesive energies for all optimized prototype structures at pressures ranging from 0 to 100 GPa calculated by both the DFT and the optimal MLP. In most cases, the MLP yields minor errors across different pressure conditions. Figure 4 also presents the distributions of (b) cohesive energy values, (c) force components, and (d) stress tensor components for dataset 2. The MLP produces narrow error distributions in all three categories, demonstrating high predictive accuracy for diverse structures and local environments, as illustrated in Fig. 2(b).
II.6 Predictive power
II.6.1 Energy-volume curves
II.6.2 Phonon properties
Figure 6 presents the phonon density of states (DOS) for diamond-, $\beta$-Sn-, SH-, Si-VI-, HCP-, and FCC-type structures at 50 GPa in elemental Si. These calculations are performed using the DFT and the two MLPs within the harmonic approximation. The finite displacement method in phonopy code [67] is employed to obtain phonon properties and thermal expansion, with supercells constructed by expanding the conventional unit cells of the respective structures. The phonon DOS calculated with the MLP constructed from dataset 1 is inconsistent with the DFT results, especially for the FCC-type structure. By comparison, the phonon DOS obtained using the MLP from dataset 2 agrees with the DFT results.
Figure 7 shows the temperature dependence of thermal expansion, calculated using the quasi-harmonic approximation, for the $\beta$-Sn-, SH-, HCP-, and FCC-type structures at 10, 30, 50, and 90 GPa, respectively. Although developing MLPs that accurately predict thermal expansion is more challenging than those designed to predict phonon DOS, the MLP developed from dataset 2 accurately predicts thermal expansion. By contrast, the MLP trained on dataset 1 produces inconsistent results and cannot compute the thermal expansion for the FCC-type structure because it predicts the FCC-type structure to be dynamically unstable.

### v2

# Global structure searches under varying temperatures and pressures using polynomial machine learning potentials: A case study on silicon

@@ I Introduction
Many elemental systems and compounds are known to exhibit polymorphism, wherein they can crystallize into multiple distinct crystal structures. Notably, numerous polymorphs have been observed, especially under high-pressure conditions. To predict such crystal structures under varying pressure conditions, ranging from ambient to high pressures, global crystal structure searches using density functional theory (DFT) calculations have proven to be a powerful and effective tool [1, 2, 3, 4]. Recent studies have also shown that machine learning potentials (MLPs) [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] significantly accelerate global structure searches by facilitating efficient energy and force calculations during local geometry optimizations [31, 32, 33, 34, 35, 36, 37, 38]. MLPs and their frameworks have been developed to perform large-scale simulations and a vast number of systematic calculations accurately and efficiently, which would be computationally expensive using DFT calculations. These MLPs represent interatomic interactions by utilizing a range of structural features that describe the neighboring atomic distribution, in conjunction with machine learning models such as artificial neural networks [5, 6, 7, 8, 9, 10, 11, 12, 13], Gaussian process models [14, 15, 16, 17, 18], and linear models [19, 20, 24, 21, 22, 23, 25]. The incorporation of MLPs facilitates robust global structure searches over extensive configurational spaces that would otherwise be computationally prohibitive using DFT calculations alone. This approach substantially reduces the risk of failing to identify the true global minimum structures.
MLPs are expected to be highly useful for performing reliable global structure searches under high-pressure conditions. At the same time, the development of MLPs capable of accurately predicting a broad range of structures, including many hypothetical configurations at elevated pressures, is essential. Although MLPs have been developed for various systems, existing models generally require modification before being applied to global structure searches. Even in the case of elemental silicon, a relatively simple system for which both conventional interatomic potentials [39, 40, 41, 42, 43] and several MLPs [16, 18, 44, 13] have been proposed, existing MLPs do not exhibit sufficient accuracy to enable global structure searches across various pressure conditions, as discussed later. Consequently, in most situations, it is necessary to develop MLPs with sufficient accuracy for high-pressure global structure prediction prior to conducting such searches.
The first part of this study presents an efficient and robust approach for developing MLPs applicable to crystal structure prediction under various pressure conditions. This approach extends an iterative procedure previously employed for global structure searches at ambient pressure using the polynomial MLP [36, 37, 38], which demonstrated high predictive accuracy across a broad range of structures and enabled the global enumeration of local minimum structures. The procedure combines random structure search (RSS) with MLP development, initiated from a dataset generated under diverse pressure conditions. As a result, the iterative process yields an MLP with high predictive power across a wide range of structures along with a set of global and local minimum structures at various pressures.
In addition to global structure searches at zero temperature, identifying globally stable structures at finite temperatures remains a significant challenge. To estimate phase stability at finite temperatures based on global structure searches, self-consistent phonon approaches [45, 46, 47, 48, 49] can be feasibly employed within practical computational limits, provided that MLPs are developed with sufficient accuracy to support such calculations. These approaches incorporate both harmonic and anharmonic vibrational effects and can predict temperature-induced dynamical stability in structures that are dynamically unstable at lower temperatures. While self-consistent phonon calculations can be readily applied to a single and simple structure, as demonstrated in previous studies [50, 51, 52, 53], their application to the broad set of local minimum structures obtained from global structure searches poses several challenges: (1) MLPs must be developed with sufficient accuracy to perform self-consistent phonon calculations across a wide range of local minimum structures and thermodynamic conditions, similar to the requirements for global structure searches. (2) Despite the acceleration offered by MLPs, the computational cost of performing self-consistent phonon calculations across many structures and conditions remains substantial. (3) These calculations are demanding particularly for low-symmetry structures, which are commonly found among local minimum structures. Consequently, a systematic and efficient procedure for performing self-consistent phonon calculations across various structures and conditions is still needed.
The second part of this study presents a procedure for systematically and accurately performing the stochastic self-consistent harmonic approximation (SSCHA) [54, 55, 56] across a wide range of structures and thermodynamic conditions, involving the development of MLPs. This procedure is applied to evaluate the free energy of local minimum structures obtained from global structure searches. Therefore, the present study also serves as a demonstration of the applicability of SSCHA calculations to complex structures that have not been previously addressed in the literature. Throughout this study, it is shown that the integrated framework, comprising the generation of DFT datasets covering diverse structures and pressure-temperature conditions, the development of accurate MLPs, robust global structure searches, and reliable SSCHA calculations, can yield a highly accurate pressure-temperature phase diagram.
The performance of the proposed procedures is evaluated in elemental Si as a prototypical example, which exhibits several experimentally observed polymorphs. Experimental studies on elemental Si have identified seven distinct phases that exist in sequence with increasing pressure: Si-I (diamond), Si-II ($\beta$-Sn), Si-XI (space group $Imma$), Si-V (simple hexagonal, SH), Si-VI (space group $Cmce$), Si-VII (hexagonal close-packed, HCP), and Si-X (face-centered cubic, FCC) within the pressure range of 0–100GPa at room temperature [57, 58, 59]. Throughout this study, we have developed a novel MLP for elemental Si, starting from a DFT dataset that was previously used to construct an earlier MLP proven to be accurate for a wide variety of structures at zero pressure, including grain boundary structures [60] and liquid configurations [61]. The current MLP also extends this accuracy to structures under pressures up to 100 GPa and significantly improves predictive performance for a broad range of local minimum structures and SSCHA calculations compared to existing MLPs.
In this study, a polynomial MLP is employed to perform the necessary calculations. This model is constructed using polynomial rotational invariants systematically derived from order parameters based on radial and spherical harmonic functions. Polynomial MLPs developed for general-purpose applications have demonstrated high predictive accuracy across a wide range of crystalline [62, 63, 64, 60] and liquid states [61]. Although the representational capacity of the potential energy surface is constrained by the use of simple polynomial functions of the invariants, compared to more flexible models such as artificial neural networks or Gaussian process models, the model estimation remains computationally efficient. This efficiency is enabled by linear regression techniques supported by optimized linear algebra libraries. Furthermore, force and stress tensor components from DFT training datasets can be straightforwardly incorporated into the fitting procedure. Even when the number of entries exceeds one million due to the inclusion of these quantities, the model coefficients can still be efficiently estimated using linear regression.
@@
An overview of the procedure proposed in this study is provided in Fig. 1, and the paper is organized as follows. Section II describes the formulation and construction procedure of the polynomial MLP. Section II.5 presents a grid search approach for obtaining an optimal MLP, which is employed in the global structure search and phase stability analysis. Section III details the iterative procedure for a reliable global structure search and presents the globally stable and metastable structures identified by the search. Section IV introduces the phase stability analysis procedure and provides the pressure-temperature phase diagram calculated using the proposed methodology. Section IV.4 assesses both the computational requirements and the predictive accuracy of the MLP for SSCHA calculations. Finally, Sec. V concludes the study.
@@ II.1 Formulation of the polynomial MLP
In addition to the model given by Eq. (S21), simpler models composed of a linear polynomial of structural features and a polynomial of a subset of the structural features are also introduced, such as
@@ II.2 Hybrid polynomial MLP models
In this study, we introduce an extended polynomial MLP model, formulated as the sum of multiple polynomial MLPs constructed using different sets of input parameters. We refer to this model as a hybrid polynomial MLP model. Such hybrid models can be implemented in a straightforward manner, particularly within the polynomial MLP framework. By incorporating both single and hybrid models, the overall model flexibility is enhanced, thereby increasing the potential for improving predictive performance. As demonstrated later, the hybrid models achieve lower prediction errors for elemental silicon compared to the single models.
The atomic energy in the hybrid polynomial MLP is expressed as the sum of two distinct MLP models, $E^{(i)}_{\rm{mdl1}}$ and $E^{(i)}_{\rm{mdl2}}$, given by
@@
Here, $D^{(i)}_{\rm{mdl1}}$ and $D^{(i)}_{\rm{mdl2}}$ represent sets of structural features generated using different input parameter sets, such as the cutoff radius, the number of radial functions, and the truncation of polynomial invariants. Specifically, these parameters include the maximum angular momentum of spherical harmonics and the polynomial order of the invariants. These hybrid polynomial MLP models extend the framework presented in Eq. (S21) and are expected to efficiently include a broader range of structural features derived from various input parameters, without requiring the consideration of intersection polynomials between the individual models.
@@ II.3 Datasets
We begin our dataset development by using an existing dataset generated from DFT calculations, comprising approximately 12,000 structures of elemental silicon. This dataset, referred to hereafter as dataset 1, was previously constructed in our earlier studies [65, 60]. Dataset 1 is derived solely from prototype structures optimized at zero pressure and therefore includes only a limited number of structures with small volumes. As a result, the MLP trained on this dataset has relatively low predictive power for various properties under high-pressure conditions, as demonstrated in Appendix B.
The procedure used to generate dataset 1 is summarized as follows. First, the atomic positions and lattice constants of 86 prototype structures [62] were optimized using DFT calculations under zero pressure. These prototypes consist of single elements with zero oxidation state from the Inorganic Crystal Structure Database (ICSD) [71], including metallic close-packed structures, covalent structures, layered structures, and those reported as high-pressure phases. Next, supercells of these optimized structures were used to introduce random lattice expansions, lattice distortions, and atomic displacements. DFT calculations were finally performed on these supercells to generate the dataset.
To develop MLPs with higher predictive accuracy under high-pressure conditions in this study, we construct an improved dataset by generating additional structures from prototype structures optimized at elevated pressures using the following procedure. First, the 86 prototype structures used to construct dataset 1 are optimized using DFT calculations at 25, 50, 75, and 100 GPa. Next, supercells of these optimized prototypes are used to create candidate structures by randomly introducing lattice expansions, distortions, and atomic displacements, following the same procedure as used for generating dataset 1. A parameter controlling the upper bounds of the magnitude for these transformations is set to 0.3 $\rm{\AA}$. Finally, to reduce uncertainty more efficiently than random selection, we apply the structure-selection procedure described in Ref. [62], extracting a total of 4,000 structures from the candidate set. The dataset comprising these structures, combined with dataset 1, will be referred to as dataset 2. When the dataset is used to estimate MLPs through regression, it is randomly split into training and test sets, with 90 % allocated to training and 10 % to testing.
@@ II.4 MLP estimation
Here, $\lambda$ represents the regularization parameter optimized to minimize prediction errors on the test set, and $\bm{W}$ is a diagonal matrix whose entries weight the energy, force, and stress tensor data. The weight settings for the energy and force entries are consistent with those proposed in Ref. [37]. In these settings, larger weights are assigned to data entries with lower energy values and smaller forces, which is critical for accurately modeling structures near local minima. The weight for the stress tensor entry, $W(\sigma_{[i],\alpha\beta})$, where $\sigma_{[i],\alpha\beta}$ denotes the stress component acting in the $\beta$ direction on a plane normal to the $\alpha$ direction in the $i$-th structure, is heuristically set as follows:
$W(\sigma_{[i],\alpha\beta})=\begin{cases}0.1\left(\dfrac{\epsilon_{\sigma}}{|\sigma_{[i],\alpha\beta}|}\right)^{0.7}\!\!&(|\sigma_{[i],\alpha\beta}|\geq\epsilon_{\sigma})\\ 0.1&(|\sigma_{[i],\alpha\beta}|<\epsilon_{\sigma})\end{cases}.$ (13)
We set the threshold $\epsilon_{\sigma}$ to 1.0 eV/cell in this study. To preserve high predictive accuracy of the regression model for the energy, force, and stress tensor data, the stress tensor entries are assigned relatively smaller weights than the energy entries.
II.5 MLP model selection
@@
As demonstrated in Ref. [62], the accuracy and computational efficiency of the polynomial MLP are trade-off properties. We therefore obtain a set of Pareto-optimal MLPs with different trade-offs between these properties through a grid search. This search covers 523 single polynomial MLP models and 640 hybrid polynomial MLP models. In the single polynomial MLP models, the cutoff radius ranges from 4 to 12 ${\rm{\AA}}$, and polynomial invariants are considered up to approximately 25,000. In the hybrid polynomial MLP models, two separate models are combined. The first model uses smaller cutoff radii (4, 5, or 6 ${\rm{\AA}}$) and 3,000–25,000 polynomial invariants, capturing a wide range of coordination environments. The second model adopts larger cutoff radii (6, 7, 8, 10, or 12 ${\rm{\AA}}$) and a varying number of polynomial invariants (1,000 to 25,000). These polynomial MLPs are developed from dataset 2 using the pypolymlp code [65, 78]. Using the prediction errors of polynomial MLPs developed with dataset 2, we select an MLP model to be used in the subsequent structure enumeration. As described later, the selected polynomial MLP is retrained using an updated dataset that combines dataset 2 with local minimum structures identified through global structure searches.
@@
A hybrid polynomial MLP model with a computational cost of 3.1 ms/atom/step is selected from the set of Pareto-optimal MLPs, balancing prediction accuracy and efficiency. The root mean square errors (RMSEs) for energy, force, and stress tensor predictions are 2.8 meV/atom, 0.055 eV/Å, and 35.9 meV/atom, respectively. To ensure practical relevance, these RMSE values are evaluated on test datasets derived from Dataset 2 that exclude structures with exceptionally high energies or large forces, which are assigned small weights. Model parameters for this selected MLP are provided in the supplemental material. The prediction errors and computational performance of other polynomial MLPs appear in Appendix A. Although some models yield slightly lower RMSEs, their overall predictive performance on key properties remains comparable to that of the selected model. Furthermore, the chosen MLP model exhibits a favorable trade-off between accuracy and computational cost. Accordingly, this model is designated as the optimal one for the present global structure searches and phase stability analyses.
Figure 3(a) compares cohesive energies for all optimized prototype structures at pressures ranging from 0 to 100 GPa calculated by both the DFT and the optimal MLP. In most cases, the MLP yields minor errors across different pressure conditions. Figure 3 also presents the distributions of (b) cohesive energy values, (c) force components, and (d) stress tensor components for dataset 2. The MLP produces narrow error distributions in all three categories, demonstrating high predictive accuracy for diverse structures and local environments, as illustrated in Fig. 2(b).
The predictive performance of the optimal MLP for the energy–volume relationship, phonon properties, and thermal expansion is presented in Appendix B. The results demonstrate that the MLP accurately reproduces these physical properties. In addition, the differences in predictive accuracy between MLPs trained on dataset 1 and dataset 2 are evaluated. These findings indicate that achieving high predictive accuracy under elevated pressures requires the inclusion of randomly generated structures derived from prototypes that are optimized at high pressures.